OpenStudy (anonymous):

How many four-digit key combinations are there if no digits are to be repeated?

OpenStudy (kropot72):

If zero is included, there are 10 digits to select from. The number of permutations of 10 digits taken 4 at a time is: \[\large 10P4 = \frac{10!}{(10-4)!}=you\ can\ calculate\]

OpenStudy (kropot72):

The right hand side of the equation can be simplified to: \[\large \frac{10!}{6!}\]

OpenStudy (kropot72):

@annatchernyak Are you there?

OpenStudy (anonymous):

@kropot72 Thank you so much ! yes its 5040

OpenStudy (kropot72):

You're welcome :)